Operations Report – March 12th

Crew 206 Operations Report 12 – MAR – 2019

SOL: 16

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 13 hours

Solar— SOC % – Turned off at 100%, Still off around 80% (1930)

Diesel Reading – 60%
Propane Reading – 75%

Ethanol Free Gasoline – 2 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 340 gallons

Water differential (static tank) – 12 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 213 gallons

Static to Loft Pump used – Yes

Water Meter: 0141745,8 gallons

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Breaking system being repared.

Hours: 53,1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Curiosity rover used: Used

Hours: 80,5 hours

Beginning charge: 100%

Ending charge: 76%

Currently charging: Yes

Spirit rover used: Used

Hours: 81,9 hours

Beginning charge: 100%

Ending charge: 44%

Currently charging: Yes

Notes on rovers: Opportunity’s brake fluid is empty, it looks like the pipeline has a leak. Atila is working on it, more informations will follow.
Curiosity showed an engine issue light during this morning’s EVA, stopping it from functioning. Restarting the rover seems to have overcome the issue for the end of the EVA.
Curiosity has been left turned on after this morning’s EVA on mistake, but it was plugged in.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not Used.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.